# Fairfax, VA, USA

### 15 Mar 2026 v1

@@ 1 Introduction
Figure 1: ASAP integration within the LVLM architecture. The plug-and-play pruning module is embedded directly within the language model backbone, operating between standard decoder layers to progressively compress the visual token sequence during the inference forward pass.
@@ 2.1 Large Vision Language Models
The remarkable capabilities of LLMs [chiang2023vicuna, touvron2023llama, bai2023qwen, touvron2023llama2, jiang2024mistral] on diverse tasks such as question-answering, reasoning, coding, etc. has inspired the extension to support multimodal data, such as images and videos, in LVLMs  [liu2023visual, bai2023qwen2, liu2024improved]. These LVLMs usually consist of an LLM backbone and a visual encoder [radford2021learning, dosovitskiy2020image, sun2023eva, zhai2023sigmoid] that embeds visual tokens before passing them to an alignment module, to align with text embeddings. More recent LVLMs [achiam2023gpt, cheng2024videollama, lin2024video, liu2024llavanext, team2024gemini, guo2024llava, bai2025qwen2, zhang2024video] capabilities have been extended to support high-resolution image and video data, resulting in a large number of visual tokens. For example, LLaVA-NeXT [liu2024llavanext] converts a high-resolution image into 4 $\times$ 4 tiles, resulting in a sequence of $\sim$ 2, 304 visual tokens. Similarly, LLaVA-OneVision [li2024llava] encodes the image into 384 $\times$ 384 tiles, producing up to 8,748 tokens, and InternVL2 [chen2024expanding] partitions an image into 448 $\times$ 448 tiles, resulting in up to 10,240 tokens. However, the massive visual token count incurs a computational burden that scales quadratically, constraining the adaptation of LVLMs. Thus, efficient inference strategies that save computational resources have become increasingly essential.
@@ 2.2 Token Pruning in LVLM/ViT
Token compression has emerged as a potential solution to the $O(N^{2})$ computational complexity of the self-attention mechanism in LVLMs. Token compression can be achieved by either pruning irrelevant tokens or merging semantically similar tokens. These strategies were first extensively studied in the context of Vision Transformers (ViTs) [cao2023pumer, chen2023diffrate, fayyaz2022adaptive, haurum2023tokens, rao2021dynamicvit, bolya2022token, feng2023efficient, kim2024token, wei2023joint] before the era of LVLMs. For example, ToMe [bolya2022token] introduced a bipartite soft matching algorithm to merge tokens based on feature similarity, while EViT [liang2022not] utilized attention scores between $[CLS]$ token and visual tokens as a proxy of token importance. In the LVLM domain, these concepts have been adapted but face new challenges. In LVLM, techniques such as LLaVA-Mini [zhang2025llava], Llavolta [chen2024efficient], and others [ye2025atp, zhang2025p] integrate pruning mechanisms but require an additional training phase. These approaches incur an additional computational burden, which undermines the primary goal of efficiency and limits their “plug-and-play” applicability. Consequently, training-free, plug-and-play approaches have received more attention recently. These methods can be categorized into several subgroups. The first group that includes LLaVA Prumerge [shang2025llava], Visionzip [yang2025visionzip], VisPruner [zhang2025beyond], and HiRED [arif2025hired] extends the ViT-era logic, relying on the attention scores between a [CLS] token and the visual tokens. However, these pruning strategies rely on the [CLS] token, which is present only in specific architectures, thereby limiting adaptability to different LVLMs. Another line of training-free work, including FastV [chen2024image], PyramidDrop [xing2024pyramiddrop], SparseVLM [zhang2024sparsevlm], and FitPrune [ye2025fit], leverages text-visual cross-attention scores to identify salient tokens. However, these techniques suffer from the critical attention sink phenomenon inherent to LVLMs [su2024roformer]. Due to positional embeddings like RoPE, later tokens receive high attention scores because of architectural functions rather than the token’s semantic content.
@@ 4 Methodology
Motivated by these structural limitations, we propose a bidirectional attention strategy to allow unrestricted information flow among visual tokens. Specifically, we relax the strict unidirectional constraint of the standard causal attention exclusively for the visual tokens, while preserving standard auto-regressive constraints for all subsequent text tokens. While recent uniform bidirectional masks [xing2024mitigating, tian2025identifying] relax these constraints, their static nature inadvertently broadcasts noise from uninformative background features. To overcome this, we propose a dynamic, targeted masking strategy to optimize visual information flow.
In the rest of this section, we present the different components of our framework. At first, we introduce the Salience-Guided Bidirectional Attention component in Section 4.1. Then, we present the Salience-Weighted Consolidation component and Budget Reallocation in Section 4.2. The ASAP framework architecture is persented in Figure 3.
@@ 4.1 Salience-Guided Bidirectional Attention
Motivated by our prior findings, we propose a Salience-Guided Bidirectional Mask (SG-BiMask) that dynamically relaxes the causal constraint for visual tokens to overcome the spatial-temporal discrepancy of strict autoregression. Rather than uniformly blinding queries to future keys, our approach introduces a target-dependent, additive penalty matrix $\mathbf{M}_{\text{bidir}}$ that scales the visibility of a future patch based on its underlying structural importance. We define $\mathbf{W}=\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d}}\in\mathbb{R}^{N\times N}$ as the pre-softmax alignment matrix computed during the standard forward pass. To extract the structural salience of each visual token without introducing auxiliary $\mathcal{O}(N^{2})$ computation, we repurpose the in-flight pre-softmax alignment matrix $\mathbf{W}$. Let $\mathcal{I}$ denote the set of indices corresponding to the visual patches. We formulate the raw attention mass $s_{j}$ of a target key token $j\in\mathcal{I}$ as its summed alignment with all visual queries $i\in\mathcal{I}$:
@@ 4.2 Salience-Weighted Consolidation and Budget Reallocation
While the ASAP mask successfully preserves high-impact visual anchors, it inherently leads to the retention of semantically redundant patches (e.g., adjacent patches of a uniform background). Directly applying standard reduction algorithms treats all features uniformly during aggregation. In the context of LVLMs, this cardinality-based averaging risks severe feature dilution, blurring the salient visual patches isolated by our bidirectional mask [bolya2022token].
To optimize the informational density of our strict $k$-token retention budget, we introduce a Salience-Weighted Consolidation mechanism. This module couples feature redundancy with the structural importance scores ($\hat{s}$) extracted during the Salience-Guided Bidirectional Masking phase.
@@ 5.1 Experimental Setup
We implemented ASAP on different LVLMs, namely LLaVA-1.5-13B [liu2023visual], and LLaVA-NeXT (LLaVA-1.6-7B) [liu2024llavanext] to demonstrate the effectiveness of our pruning recipe across different models. LLaVA-1.5-13B encodes a standard 336x336 image using 576 visual tokens, supporting low-resolution images. On the other hand, LLaVA NeXT supports high-resolution images by partitioning an image into $4\times 4$ tiles, generating up to 2880 tokens. This 4x increase in token count over LLaVA-1.5 presents a significant computational bottleneck, making it an ideal testbed for our pruning framework.
@@ 5.2 Main Results
We evaluated ASAP against leading training-free baselines on diverse image understanding tasks. Our evaluation was conducted at two distinct computational budgets, corresponding to $\sim$32.99% and $\sim$21.89% of the vanilla LLaVA-1.5-7b’s total FLOPs. To normalize comparisons across all benchmarks, we report an average performance metric, where scores are scaled relative to the vanilla model.
ASAP establishes impressive results in training-free pruning, demonstrating a superior trade-off between computational cost and model performance. As detailed in Table 1, our method is highly effective across diverse pruning budgets. At a high-efficiency setting (retaining 128 tokens at a 20.90% FLOPs ratio), our method retains 97.10% of the vanilla model’s average performance. Under this extreme compression, ASAP sees minimal degradation on benchmarks such as MMBench (66.18 vs. 67.71) and MMBench-CN (62.46 vs. 63.60) while robustly maintaining performance on VQA-v2 (76.26) and GQA (58.76).
@@
Compared to existing methods, ASAP consistently outperforms the strongest prior baseline, Divprune. ASAP beats Divprune’s average performance retention by 1.08% at the 128-token budget (97.10% vs. 96.02%) and by 2.00% at the 192-token budget (98.68% vs. 96.68%), confirming its definitive superiority in optimizing the efficiency-accuracy frontier.
Table 1: Performance comparison of ASAP with other pruning strategies on LLaVA-1.5-13B. Ratio represents the FLOPs ratio (percentage of vanilla FLOPs retained). Avg refers to the average percentage performance across all benchmarks compared to the vanilla model. Best scores are highlighted in red.
@@ 5.3 Results on High-Resolution Images
To validate the generalizability of ASAP, we integrated it into LLaVA-NeXT, a model designed for high-resolution vision tasks. LLaVA-NeXT partitions input images into $4\times 4$ tiles, resulting in a large and computationally expensive sequence of up to $\sim 2,880$ visual tokens. While this tiling approach enhances performance on fine-grained tasks, it also introduces significant token redundancy, making it an ideal testbed for our method. We evaluated ASAP on LLaVA-NeXT using two distinct retention budgets of 33.33% and 22.23%, as previously. As shown in Table 2, at the 33.33% budget ($\sim$960 tokens), our method achieves performance that is comparable to, and in some cases surpasses, the vanilla LLaVA-NeXT model. Moreover, with a more aggressive pruning rate retaining only 22.23% ($\sim$640 tokens) of the original tokens, ASAP’s performance degrades by only $\sim$0.98 points on average, demonstrating robust high-compression performance. Furthermore, ASAP establishes a new state-of-the-art, outperforming the VisPruner baseline by 1.80% and 2.32% at the 33.33% and 22.23% retention budgets, respectively.
Table 2: Performance comparison of ASAP with other pruning strategies on LLaVA-NeXT-7B. Avg refers to the average percentage performance across all benchmarks compared to the vanilla model. Best scores are highlighted in red.
@@ 5.4 Results on Different LVLM Architecture
To further validate the generalizability of ASAP, we integrated it into InternVL 2.5 [chen2024expanding], a modern model architecture designed for high-resolution vision tasks. Similar to other high-resolution architectures, InternVL employs a dynamic tiling strategy that generates a massive sequence of visual tokens. While this enhances performance on fine-grained tasks, it also introduces significant token redundancy, making it an ideal testbed for our method. As shown in Table 3, with the FLOPS ratio of 31.56% ($\sim$ 510 tokens), ASAP maintains the model’s performance, experiencing a minimal average degradation of only 1.24 points compared to the vanilla model while drastically reducing the computational cost from 12.575 to 3.969 FLOPs. Moreover, ASAP remains robust under a more aggressive pruning rate of 21.17% FLOPs ratio ($\sim$ 340 tokens) by minimizing the accuracy drop. Furthermore, ASAP establishes superior efficiency over existing approaches, outperforming the FastV [chen2024image] baseline across average benchmark scores by 1.59% and 0.67% at the 31.56% and 21.17% retention budgets, respectively. These performances establish the generalizability of the ASAP pruning framework to modern LVLM architectures.
@@ 5.5 Ablation Studies
Figure 4: Ablation of ASAP modules on TextVQA and MME. FastV prunes naively using Top-k selection. Important uses Salience-Guided Bidirectional attention values. ASAP adds Salience-Weighted Consolidation and Budget Reallocation. The progressive gains confirm that combining attention-driven selection with feature consolidation maximizes semantic density.
Bidirectional (Module 1): Next, we replace the standard causal attention with our Salience-Guided Bidirectional attention values for token selection. This provides a significant performance leap, yielding an absolute increase of +3.23 on TextVQA (to 56.28) and +92.50 on MME (to 1375.31) over the FastV baseline. This validates the hypothesis that relaxing strict causal constraints and permitting bidirectional information flow allows the model to better identify and retain structurally critical visual anchors.
@@ 5.6 Efficiency analysis
To evaluate the system efficiency of our approach, we perform a comparative analysis of FLOPs, KV cache storage, GPU memory, and CUDA latency alongside task accuracy against FastV on the LLaVA-1.5-7B [liu2023visual] architecture. As shown in Table 4, reducing the visual token count to 128 yields a substantial decrease in computational cost, dropping the required operations from 3.817 to 0.836 TFLOPs. Under equivalent token retention budgets, our method consistently achieves higher accuracy than the FastV baseline while simultaneously reducing the KV cache storage footprint. For instance, at the 128-token budget, our approach improves the aggregate accuracy score by over 52 points while requiring approximately 15% less storage space than FastV. Although our dynamic token selection mechanism introduces a marginal inference overhead of roughly two milliseconds, it provides a strictly superior balance between memory efficiency and downstream model performance.
@@
Method # Token FLOPs (T) Accuracy Storage (MB) GPU Memory (GB) CUDA Time (ms)

### 17 Mar 2026 v2

@@ 1 Introduction
Figure 1: ASAP integration within the LVLM architecture. The plug-and-play pruning module is embedded directly within the language model backbone, operating between standard decoder layers to compress the visual token sequence during the inference forward pass.
@@ 2.1 Large Vision Language Models
The remarkable capabilities of LLMs [chiang2023vicuna, touvron2023llama, bai2023qwen, touvron2023llama2, jiang2024mistral] on diverse tasks such as question-answering, reasoning, coding, etc. has inspired the extension to support multimodal data, such as images and videos, in LVLMs [liu2023visual, bai2023qwen2, liu2024improved]. These LVLMs usually consist of an LLM backbone and a visual encoder [radford2021learning, dosovitskiy2020image, sun2023eva, zhai2023sigmoid] that embeds visual tokens before passing them to an alignment module, to align with text embeddings. The capabilities of more recent LVLMs [achiam2023gpt, cheng2024videollama, lin2024video, liu2024llavanext, team2024gemini, guo2024llava, bai2025qwen2, zhang2024video] have been extended to support high-resolution image and video data, resulting in a large number of visual tokens. For example, LLaVA-NeXT [liu2024llavanext] converts a high-resolution image into 4 $\times$ 4 tiles, resulting in a sequence of $\sim$ 2, 304 visual tokens. Similarly, LLaVA-OneVision [li2024llava] encodes the image into 384 $\times$ 384 tiles, producing up to 8,748 tokens, and InternVL2 [chen2024expanding] partitions an image into 448 $\times$ 448 tiles, resulting in up to 10,240 tokens. However, the massive visual token count incurs a computational burden that scales quadratically, constraining the adaptation of LVLMs. Thus, efficient inference strategies that save computational resources have become increasingly essential.
@@ 2.2 Token Pruning in LVLM/ViT
Token compression has emerged as a potential solution to the $O(N^{2})$ computational complexity of the self-attention mechanism in LVLMs. Token compression can be achieved by either pruning irrelevant tokens or merging semantically similar tokens. These strategies were first extensively studied in the context of Vision Transformers (ViTs) [cao2023pumer, chen2023diffrate, fayyaz2022adaptive, haurum2023tokens, rao2021dynamicvit, bolya2022token, feng2023efficient, kim2024token, wei2023joint] before the era of LVLMs. For example, ToMe [bolya2022token] introduced a bipartite soft matching algorithm to merge tokens based on feature similarity, while EViT [liang2022not] utilized attention scores between $[CLS]$ token and visual tokens as a proxy of token importance. In the LVLM domain, these concepts have been adapted but face new challenges. In LVLM, techniques such as LLaVA-Mini [zhang2025llava], Llavolta [chen2024efficient], and others [ye2025atp, zhang2025p] integrate pruning mechanisms but require an additional training phase. These approaches incur an additional computational burden, which undermines the primary goal of efficiency and limits their “plug-and-play” applicability. Consequently, training-free, plug-and-play approaches have received more attention recently. These methods can be categorized into several subgroups. The first group that includes LLaVA Prumerge [shang2025llava], Visionzip [yang2025visionzip], VisPruner [zhang2025beyond], and HiRED [arif2025hired] extends the ViT-era logic, relying on the attention scores between a [CLS] token and the visual tokens. However, these pruning strategies rely on the [CLS] token, which is present only in specific architectures, thereby limiting adaptability to different LVLMs. Another line of training-free work, including FastV [chen2024image], PyramidDrop [xing2024pyramiddrop], SparseVLM [zhang2024sparsevlm], and FitPrune [ye2025fit], leverages text-visual cross-attention scores to identify salient tokens. However, these techniques suffer from the critical attention shift phenomenon inherent to LVLMs [su2024roformer]. Due to positional embeddings like RoPE, later tokens receive high attention scores because of architectural functions rather than the token’s semantic content. Our work structurally addresses this limitation, enabling an efficient pruning recipe that rectifies naive attention-based selection.
@@ 4 Methodology
Motivated by the observed structural limitations in Section 3, we propose a bidirectional attention strategy to allow unrestricted information flow among visual tokens. Specifically, we relax the strict unidirectional constraint of the standard causal attention exclusively for the visual tokens, while preserving standard auto-regressive constraints for all subsequent text tokens. While recent uniform bidirectional masks [xing2024mitigating, tian2025identifying] relax these constraints, their static nature inadvertently broadcasts noise from uninformative background features. To overcome this, we propose a dynamic, targeted masking strategy to optimize visual information flow.
@@ 4.1 Salience-Guided Bidirectional Attention
Motivated by our prior findings in Section 3, we propose a Salience-Guided Bidirectional Mask (SG-BiMask) that dynamically relaxes the causal constraint for visual tokens to overcome the spatial-temporal discrepancy of strict autoregression. Rather than uniformly blinding queries to future keys, our approach introduces a target-dependent, additive penalty matrix $\mathbf{M}_{\text{bidir}}$ that scales the visibility of a future patch based on its underlying structural importance. We define $\mathbf{W}=\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d}}\in\mathbb{R}^{N\times N}$ as the pre-softmax alignment matrix computed during the standard forward pass. To extract the structural salience of each visual token without introducing auxiliary $\mathcal{O}(N^{2})$ computation, we repurpose the in-flight pre-softmax alignment matrix $\mathbf{W}$. Let $\mathcal{I}$ denote the set of indices corresponding to the visual patches. We formulate the raw attention mass $s_{j}$ of a target key token $j\in\mathcal{I}$ as its summed alignment with all visual queries $i\in\mathcal{I}$:
@@ 4.2 Salience-Weighted Consolidation and Budget Reallocation
While the ASAP mask successfully preserves high-impact visual anchors, it inherently leads to the retention of semantically redundant patches (e.g., adjacent patches of a uniform background). Directly applying standard reduction algorithms treats all features uniformly during aggregation [bolya2022token]. In the context of LVLMs, this cardinality-based averaging risks severe feature dilution, blurring the salient visual patches isolated by our bidirectional mask [wen-etal-2025-stop].
To optimize the informational density of our predefined token retention budget (defined in Section 5.25.3, and 5.4, respectively), we introduce a Salience-Weighted Consolidation mechanism. This module couples feature redundancy with the structural importance scores ($\hat{s}$) extracted during the Salience-Guided Bidirectional Masking phase in Section 4.1.
@@ 5.1 Experimental Setup
We implemented ASAP on different LVLMs, namely LLaVA-1.5-13B [liu2023visual], and LLaVA-NeXT (LLaVA-1.6-7B) [liu2024llavanext] to demonstrate the effectiveness of our pruning recipe across different models. LLaVA-1.5-13B encodes a standard $336\times 336$ image using 576 visual tokens, supporting low-resolution images. On the other hand, LLaVA NeXT supports high-resolution images by partitioning an image into $4\times 4$ tiles, generating up to 2,880 tokens. This $4\times$ increase in token count over LLaVA-1.5 presents a significant computational bottleneck, making it an ideal testbed for our pruning framework. In all our experiments, we implement ASAP once at the layer $k=2$ of the corresponding backbone LLM.
@@ 5.2 Main Results
We evaluated ASAP against leading training-free baselines on diverse image understanding tasks. Our evaluation was conducted at two distinct computational budgets, corresponding to $\sim$32.99% and $\sim$21.89% of the vanilla LLaVA-1.5-7b’s total FLOPs. The procedure to obtain the theoretical FLOPs is outlined in Appendix 0.B. We selected these budgets to assess ASAP’s efficiency-to-accuracy tradeoff at moderate and aggressive pruning levels, respectively. To normalize comparisons across all benchmarks, we report an average performance metric, where scores are scaled relative to the vanilla model.
ASAP establishes impressive results in training-free pruning, demonstrating a superior trade-off between computational cost and model performance. As detailed in Table 1, our method is highly effective across diverse pruning budgets. At a high-efficiency setting (retaining 128 tokens at a 20.90% FLOPs ratio), our method retains 97.10% of the vanilla model’s average performance. Under this extreme compression, ASAP sees minimal degradation on benchmarks such as MMBench (66.18 vs. 67.71) and MMBench-CN (62.46 vs. 63.60) while robustly maintaining performance on TextVQA (58.88 vs 61.19).
@@
Compared to existing methods, ASAP consistently outperforms the strongest prior baseline, DivPrune [alvar2025divprune]. ASAP beats DivPrune’s average performance retention by 1.08% at the 128-token budget (97.10% vs. 96.02%) and by 2.00% at the 192-token budget (98.68% vs. 96.68%), confirming its definitive superiority in optimizing the efficiency-to-accuracy frontier.
Table 1: Performance comparison of ASAP with other pruning strategies on LLaVA-1.5-13B [liu2023visual]. Ratio represents the FLOPs ratio (percentage of vanilla FLOPs retained). Avg refers to the average percentage performance across all benchmarks compared to the vanilla model. Best scores are highlighted in red.
@@ 5.3 Results on High-Resolution Images
To validate the generalizability of ASAP, we integrated it into LLaVA-NeXT, a model designed for high-resolution vision tasks. LLaVA-NeXT partitions input images into $4\times 4$ tiles, resulting in a large and computationally expensive sequence of up to $\sim 2,880$ visual tokens. While this tiling approach enhances performance on fine-grained tasks, it introduces significant token redundancy, making it an ideal testbed for our method. We evaluated ASAP on LLaVA-NeXT using two distinct retention budgets of 33.33% and 22.23%, as previously. In addition to FastV, we use Vispruner [zhang2025beyond] as a baseline over here since its LLaVA NeXT implementation is available publicly. As shown in Table 2, at the 33.33% budget ($\sim$960 tokens), our method achieves performance that is comparable to, and in some cases surpasses, the vanilla LLaVA-NeXT model. Moreover, with a more aggressive pruning rate retaining only 22.23% ($\sim$640 tokens) of the original tokens, ASAP’s performance degrades by only $\sim$0.98 points on average, demonstrating robust high-compression performance. Furthermore, ASAP establishes a new state-of-the-art, outperforming the VisPruner [zhang2025beyond] baseline by 1.80% and 2.32% at the 33.33% and 22.23% retention budgets, respectively.
Table 2: Performance comparison of ASAP with other pruning strategies on LLaVA-NeXT-7B [liu2024llavanext]. Avg refers to the average percentage performance across all benchmarks compared to the vanilla model. Best scores are highlighted in red.
@@ 5.4 Results on Different LVLM Architecture
To further validate the generalizability of ASAP, we integrated it into InternVL 2.5 [chen2024expanding], a modern model architecture designed for high-resolution vision tasks. Similar to other high-resolution architectures, InternVL employs a dynamic tiling strategy that generates a massive sequence of visual tokens. While this enhances performance on fine-grained tasks, it also introduces significant token redundancy, making it an ideal testbed for our method. Since other baselines’s InternVL 2.5 implementation is not available publicly, we use FastV as the baseline over here. As shown in Table 3, with the FLOPS ratio of 31.56% ($\sim$ 510 tokens), ASAP maintains the model’s performance, experiencing a minimal average degradation of only 1.24 points compared to the vanilla model while drastically reducing the computational cost from 12.575 to 3.969 FLOPs. Moreover, ASAP remains robust under a more aggressive pruning rate of 21.17% FLOPs ratio ($\sim$ 340 tokens) by minimizing the accuracy drop. Furthermore, ASAP establishes superior efficiency over existing approaches, outperforming the FastV [chen2024image] baseline across average benchmark scores by 1.59% and 0.67% at the 31.56% and 21.17% retention budgets, respectively. These performances establish the generalizability of the ASAP pruning framework to modern LVLM architectures.
Additionally, ASAP’s SG-BiMask component is orthogonal to the existing attention-based pruning methods and can be integrated to increase their pruning effectiveness. The increase in accuracy after integrating SG-BiMask to attention-based pruning is demonstrated in Appendix 0.C.1.
@@ 5.5 Ablation Study
Figure 4: Ablation of ASAP modules on TextVQA and MME. FastV prunes naively using Top-$k$ selection. Bidirectional uses Salience-Guided Bidirectional attention values. ASAP adds Salience-Weighted Consolidation and Budget Reallocation. The progressive gains confirm that combining attention-driven selection with feature consolidation maximizes semantic density.
Bidirectional (Module 1): Next, we replace the standard causal attention with our Salience-Guided Bidirectional attention values for token selection. This provides a significant performance leap, yielding an absolute increase of +3.23 on TextVQA (to 56.28) and +92.50 on MME (to 1375.31) over FastV. This validates the hypothesis that relaxing strict causal constraints and permitting bidirectional information flow allow the model to better identify and retain structurally critical visual anchors.
@@ 5.6 Efficiency Analysis
To evaluate the system efficiency of our approach, we perform a comparative analysis of FLOPs, KV cache storage, GPU memory, and CUDA latency alongside task accuracy against FastV [chen2024image] on the LLaVA-1.5-7B [liu2023visual] architecture. As shown in Table 4, reducing the visual token count to 128 yields a substantial decrease in computational cost, dropping the required operations from 3.817 to 0.836 TFLOPs. Under equivalent token retention budgets, our method consistently achieves higher accuracy than FastV while simultaneously reducing the KV cache storage footprint. For instance, at the 128-token budget, our approach improves the aggregate accuracy score by over 52 points while requiring approximately 15% less storage space than FastV. Although our dynamic token selection mechanism introduces a marginal inference overhead of roughly two milliseconds, it provides a strictly superior balance between memory efficiency and downstream model performance.
@@
Method # Token FLOPs (T) Percption score Storage (MB) GPU Memory (GB) CUDA Time (ms)